OpenStudy (hpfan101):

Differentiate: \[\frac{ x^2e^x }{ x^2+e^x }\]

OpenStudy (empty):

Are you familiar with the quotient rule?

OpenStudy (unklerhaukus):

\[\large\left(\frac uv\right)' = \frac{u'v-uv'}{v^2}\]

OpenStudy (unklerhaukus):

@hpfan101, can you identify \(u\) and \(v\)? can you find the derivatives \(u'\) and \(v'\)?

OpenStudy (hpfan101):

I'm somewhat familiar with the quotient rule.

OpenStudy (hpfan101):

Would it be: \[\frac{ (x^2e^x)'(x^2+e^x)-(x^2e^x)(x^2+e^x)' }{ (x^2+e^x)^2 }\]

OpenStudy (unklerhaukus):

yeah, that's right, now simplify those derivatives

OpenStudy (hpfan101):

Okay, what I'm not sure about is the first set we would find the derivative of. Would the derivative of \[(x^2e^x)'\] be \[2xe^x\]?

OpenStudy (unklerhaukus):

use the product rule \[(fg)' = f'g+fg'\]

OpenStudy (hpfan101):

Oh okay. So it would be: \[(x^2)'(e^x)+(x^2)(e^x)'= 2x(e^x)+x^2(e^x) =2xe^x+x^2e^x\]

OpenStudy (unklerhaukus):

that's right and if you factor it \(= (2x+x^2)e^x\)

OpenStudy (hpfan101):

Oh okay! Thank you! I think I can figure out the rest.

OpenStudy (unklerhaukus):

to compute\[(x^2+e^x)'\] use the sum rule \[(a+b)'=a'+b'\]

OpenStudy (hpfan101):

Okay, will do! Thanks!

OpenStudy (unklerhaukus):

NB: expanding the square in the denominator will not simplify your result (in this case)

OpenStudy (hpfan101):

Ah, okay. I see. It'd be easier to leave the denominator how it is.